OpenStudy (liizzyliizz):

if f(x)= ae^kx and f'(x)/f(x)= -5/2 then k= Help lol (Relearning calculus and basic rules :P )

OpenStudy (bahrom7893):

find the derivative of f(x): f'(x) = a*e^(kx)*k (Chain rule)

OpenStudy (bahrom7893):

f'(x) = a*k*e^(kx) f'(x)/f(x) = -5/2, so a*k*e^(kx)/[a*e^(kx)] = -5/2

OpenStudy (anonymous):

\[f(x)=ae ^{kx}\]\[f'(x)=ake ^{kx}\]reaplce:\[\frac{ake ^{kx}}{ae ^{kx}}=-\frac{5}{2}\]\[k=-\frac{5}{2}\]

OpenStudy (bahrom7893):

a*e^(kx) on top and bottom cancel out, so u're left with k = -5/2

OpenStudy (anonymous):

remember: \[f(x)=e ^{u} \]\[f \prime(x)=e ^{u}u \prime\]

OpenStudy (liizzyliizz):

O.o This is so simple lol I am embarrassed I got this wrong now :(

OpenStudy (anonymous):

dont forget, click good answer :D jeje

OpenStudy (liizzyliizz):

lol there you go :) Thanks for the help.